OpenStudy (swissgirl):

Find the coefficient of \( a^3b^{10} \) in the expansion of \( (a+b)^{13} \)

OpenStudy (swissgirl):

@SmoothMath can u help me

OpenStudy (anonymous):

Hint: Pascal's Muthaflutterin' triangle. Google that ish.

OpenStudy (swissgirl):

lol

OpenStudy (anonymous):

Here n= 13 you want coefficient of : \(a^3b^{10}\) So find: \[\large ^{13}C_{10}\] This will be the coefficient..

OpenStudy (swissgirl):

I am getting 286 which isnt a number on pascals triangle

OpenStudy (anonymous):

\[\large ^{13}C_{10} = \frac{13!}{(13-10)! \times 10!} = \frac{13 \times 12 \times 11}{3 !} \implies 13 \times 2 \times 11\]

OpenStudy (swissgirl):

but check on pascals triangle there is no 286 by 13

OpenStudy (anonymous):

Okay I will check..

OpenStudy (anonymous):

See: According to me it is: \[\\large \^\{\1\3\}C_\{\1\0\}\\]

OpenStudy (swissgirl):

ya that is what i thought tooo

OpenStudy (swissgirl):

U know what ill sleep over it and continue it tom. Its nearly 4 am and i gotta be up early. A new day bring new surprises

OpenStudy (anonymous):

do you have answer with you for this question??

OpenStudy (swissgirl):

nope i dont

OpenStudy (swissgirl):

Ill just sit over it tom

OpenStudy (anonymous):

Wait I tell you all the values of triangle for n = 13..

OpenStudy (swissgirl):

yaaaaa like i am assuming that it is 186

OpenStudy (anonymous):

No..

OpenStudy (swissgirl):

thats the one that makes most sense

OpenStudy (anonymous):

no I got the answer.. 13th row contains 286..

OpenStudy (swissgirl):

hmmmm

OpenStudy (swissgirl):

oh ya it does i guess its a mistake on that website

OpenStudy (swissgirl):

alrighty thanks guys Gnite

OpenStudy (anonymous):

Yes there is a mistake.. Good Night @swissgirl

OpenStudy (swissgirl):

ohhhhhhh ok Thankss

OpenStudy (anonymous):

Welcome dear..

OpenStudy (anonymous):

Sorry there will come 286 in place of 186 in the link you provided.. last one is 13th row there..

OpenStudy (mimi_x3):

An easier method that I would use than the pascals triangle is the Binomial Expansion: \((a+b)^{n} = \binom{n}{r} *(a)^{n-r}b^{r}\) \[\binom{13}{r} *(a)^{13-r}*b^{r} \] \[\text{So you are looking for} ~a^{3}b^{10} then ~r = 10 \] \[\binom{13}{10} *(a)^{13-10}*b^{10} => 286a^3b^{10}\] Well this is an easier method to me than using the Pascals traingle which is confusing

OpenStudy (mimi_x3):

And the answer is correct; you can check it here: http://www.wolframalpha.com/input/?i=%28a%2Bb%29^13